# Evaluation and Facilitation of Online Discussions in the LLM Era: A Survey

### v1

@@ Abstract
We present a survey of methods for assessing and enhancing the quality of online discussions, focusing on the potential of Large Language Models (LLMs). While online discourses aim, at least in theory, to foster mutual understanding, they often devolve into harmful exchanges, such as hate speech, threatening social cohesion and democratic values. Recent advancements in LLMs enable facilitation agents that not only moderate content, but also actively improve the quality of interactions. Our survey synthesizes ideas from Natural Language Processing (NLP) and Social Sciences to provide (a) a new taxonomy on discussion quality evaluation, (b) an overview of intervention and facilitation strategies, along with a new taxonomy on conversation facilitation datasets, (c) an LLM-oriented roadmap of good practices and future research directions, from technological and societal perspectives.
@@ 1 Introduction
Social Science provides theories and applications for the facilitation of a discussion, but in specific contexts, such as teaching/learning Mansour (2024) or clinical discussions Gelula (1997), without much research conducted for thread-like discussions.
Improving the quality of online discussions presupposes being able to define and measure discussion quality. Here, work from Social Science, but also Argument Mining (AM) Lapesa et al. (2024), can again provide several ideas on dimensions (aspects) of discussion quality, such as logical cohesion and constructiveness, as well as ideas on methods to measure quality along each dimension.
This paper surveys research from Social Science and relevant NLP areas (e.g., AM, Sentiment Analysis, Toxicity Detection), focusing on how Large Language Models can facilitate human discussions—similar to human facilitators Kahane (2013). While prior studies have explored LLM-facilitated discussions Burton et al. (2024); Aher et al. (2023); Beck et al. (2024); Schroeder et al. (2024); Small et al. (2023); Cho et al. (2024), their connection to Social Science remains underexplored. In this survey, we include methods from Social Science (e.g., linguistics) when discussing approaches for evaluating online discussions, as well as when exploring intervention strategies (e.g., facilitative tactics). LLM-based facilitation again presupposes defining and evaluating discussion quality. This is even more necessary in the case of LLMs, because of their rapid deployment, potential biases, and long-term societal consequences. Without continuous and thorough assessment, we risk implementing LLM-based facilitation systems that may be ineffective, biased, or even harmful, before their full implications are understood.
Therefore, we survey discussion evaluation aspects and their feasibility with LLMs, introducing a new taxonomy inspired by deliberation studies. We map tasks suited for ML models, LLMs, and humans, aggregate multidimensional insights on facilitation strategies, and outline future capabilities of LLMs. Additionally, we compare major datasets and introduce a taxonomy of tasks relevant to facilitation. Our work focuses mostly on written thread-like discussions (e.g., Reddit threads).
Our findings show that (a) many discussion evaluation aspects coexist in literature; we propose a precise taxonomy to alleviate this, (b) advancements in machine and machine-enhanced human intervention show significant promise in improving the quality and timeliness of facilitation methods; (c) there is a general lack of datasets for studying facilitation; we aggregate existing datasets. We posit that LLM-generated discussions, could become an asset to test facilitation strategies in diverse discussions.
@@ 2 Terminology
Given the numerous aspects to consider regarding discussion quality and facilitation, we narrow the scope of the survey and clarify the terminology we use. We highly recommend that the reader refer to the Terminology Section of Appendix C and, especially, Table 3, where we explain our findings with regard to the terms used in the literature.
@@
We find that the terms ‘moderation’ and ‘facilitation’ usually refer to the same process in the literature, with ‘facilitation’ being used more in Social Science, and ‘moderation’ in NLP Vecchi et al. (2021). Thus, we will be using them interchangeably in this survey.
@@
This survey mainly focuses on ‘Real-Time, Ex-Post-moderation’, i.e., moderation happening after the user has posted some content. This is different from pre-moderation approaches, such as nudging users before they post harmful content Argyle et al. (2023), or delaying the posting of user content until a moderator/facilitator has had the chance to check it.
@@
The definitions of these terms often vary across literature Russmann and Lane (2016); Goñi (2024). We focus on discussions, a general term for verbal/written exchanges Russmann and Lane (2016), and deliberations, a term for structured discussions focusing on opinion sharing Degeling et al. (2015); Lo and McAvoy (2023). This is in contrast to the collaborative nature of dialogues Rose-Redwood et al. (2018); Bawden (2021); Goñi (2024) and the competitive nature of debates Lo and McAvoy (2023).
@@
(or “threads”) Seering (2020) are discussions which start from an Original Post (OP) with subsequent comments replying to either the OP or to other comments.
@@ 3 Comparison to Other Surveys
Few surveys have considered discussion quality evaluation and facilitation from a perspective that encompasses ideas from both Social Science and Computer Science, as the present survey does. The closest related surveys are those by Wachsmuth et al. (2024) and Vecchi et al. (2021). Wachsmuth et al. (2024) focus primarily on discussion evaluation on its own, rather than exploring how it relates to facilitation, which is one of the main goals of our survey. Furthermore, their survey predominantly centers on concepts from AM. On the other hand, the survey of Vecchi et al. (2021) is rooted largely in AM research. They point out that traditional AM has prioritized the logical structure and soundness of arguments, while overlooking other important dimensions, such as civility, respectfulness, inclusiveness, originality, and the broader impacts of discussions—such as encouraging mutual understanding, convergence, and problem-solving. In other words, they argue that advancing AM for social good requires a collaborative effort between AM and Social Science. Building on this notion, our survey focuses on both discussion evaluation and facilitation, incorporating ideas from Social Science into NLP-based approaches.
4 Survey Methodology
The search and article selection of this survey was conducted using specific keywords in academic search engines (e.g., Google Scholar, Semantic Scholar, Scopus), digital libraries and repositories (e.g., ACL Anthology, ACM Digital Library, IEEE Xplore, JSTOR). We focused on peer-reviewed publications written in English between 2014 and 2024, granting exceptions only for established works predating this period. Additionally, we reviewed other cited papers that appeared highly relevant, provided they were peer-reviewed and cited by more than 20 citations of other researchers, unless the topic is very niche, in which case we judge by its content. The search strategy incorporated keywords and phrases related to LLMs, discussion facilitation, and discussion evaluation. The list of keywords used is provided in Appendix B in Table 2. The search was further informed by existing survey articles, such as those by Vecchi et al. (2021) and Wachsmuth et al. (2024), which served as starting points both for identifying relevant literature and for specifying the vocabulary used in the keyword search.
@@ 5 Discussion Quality Evaluation
According to Kies (2022), deliberation quality is assessed along three dimensions: (1) Deliberative Presence, examining inclusion (ability of all interested parties to participate in the deliberation process), discursive equality (equal opportunities of articulation), justification (well-reasoned arguments), and reciprocity (participants listen and react to each other); (2) Deliberative Attitudes, covering reflexivity (being open-minded to be convinced by the arguments of others), empathy (shared understanding towards other views and opinions), and sincerity (disclose true positions and beliefs); (3) Deliberative Outcomes, assessing plurality (range of opinions) and external impact on broader political discourse. However, the focus of this survey is online written discussions that do not focus exclusively on the deliberative goal of discourse Gerber et al. (2018), unlike the work of Kies (2022).
For that reason, we also take into account Social Science, which offers different theories and definitions about Argument Quality (AQ), and which focuses not only on the argument itself but also on the interaction and social dynamics between participants Falk and Lapesa (2023). In particular, we leverage Social Science (particularly Deliberative Theory (DT)) quality notions Bächtiger et al. (2022, 2010); Steenbergen et al. (2003), and discussion quality aspects Falk and Lapesa (2023). We expand the work of Kies (2022), defining a new broader taxonomy for the employed discussion quality aspects.
Figure 1: Proposed taxonomy of evaluation aspects.
@@ 5.1 Structure and Logic
AQ is a multidimensional concept assessed through logical, rhetorical, and dialectical aspects Wachsmuth et al. (2017). Logical quality evaluates coherence via premise believability, relevance, and sufficiency. Rhetorical quality incorporates credibility, emotional appeal, and clarity Ziegenbein et al. (2023); Ivanova et al. (2024). Dialectical quality measures discourse engagement and argument robustness Wachsmuth et al. (2017). Empirical studies leverage NLP and ML, and recently LLMs, to automate AQ assessment Ziems et al. (2024). LLMs demonstrate strong annotation capabilities, performing comparably to human evaluators Mirzakhmedova et al. (2024); Rescala et al. (2024), excelling in comparative argument evaluation Wang et al. (2023), AM, and synthesis Chen et al. (2024); Irani et al. (2024); Anastasiou and De Liddo (2024). For a broader review, see the work of Wachsmuth et al. (2024) and Lauscher et al. (2022).
@@
Coherence evaluates logical consistency, while flow assesses smooth progression in discussions Li et al. (2021). Initial efforts in coherence assessment focused on dialogue systems using language features Zhang et al. (2018a), discourse structure Barzilay and Lapata (2008), and graph-based representations Huang et al. (2020); Zhang et al. (2021). To assess both aspects, LLMs are increasingly used for coherence evaluation at the comment or whole discussion level Zhang et al. (2024), with proprietary models (e.g., GPT-4) excelling, and fine-tuned open-source models showing promise Mendonca et al. (2024); Zhang et al. (2023).
@@
Turn-taking patterns (e.g., how often participants speak in a discussion, to which other speakers they tend to reply or not) inform coherence evaluation Cervone and Riccardi (2020), constructiveness prediction Niculae and Danescu-Niculescu-Mizil (2016), and facilitation analysis Schroeder et al. (2024). Studies have used entropy Niculae and Danescu-Niculescu-Mizil (2016), Gini coefficients Schroeder et al. (2024), and neural architectures for modeling turn-taking dynamics Chang and Danescu-Niculescu-Mizil (2019); Li et al. (2021), as well as turn-taking visualization tools El-Assady et al. (2017); Hoque and Carenini (2016).
Language Features
Language features have been used to help model content and expression in online discussions Wilson et al. (1984). Early methods used lexicons for sentiment, toxicity, politeness and collaboration evaluations, aspects closely related to discussion quality Lawrence et al. (2017); Avalle et al. (2024). Deep learning models leverage word embeddings instead De Kock and Vlachos (2021).
@@
Rooted in Speech Act Theory Austin (1975); Searle (1969), dialogue acts characterize dialogue turns (e.g., disagreement, elaboration) to analyze interaction dynamics Ferschke et al. (2012). Various taxonomies exist, including generic Stolcke et al. (2000) and domain-specific frameworks Zhang et al. (2017); Al-Khatib et al. (2018). LLMs, in this case, can serve as dialogue act annotators Ziems et al. (2024); Cimino et al. (2024); Martinenghi et al. (2024), aiding in the distinction between effective and ineffective discussions based on the patterns of interaction and the communicative strategies and goals of the participants Ziems et al. (2024).
@@
Pragmatic comprehension (how context affects meaning) is essential in human communication, where the intended meaning often differs from the verbalized expression (e.g., in implicature). Humans resolve such ambiguities using social and common-sense knowledge. NLP models need to be evaluated on their pragmatic understanding Al-Khatib et al. (2018). Research shows that LLM-fine-tuning enhances implicature comprehension Ruis et al. (2023), with GPT-4 achieving human-level performance through chain-of-thought prompting. While LLMs perform well in some pragmatic tasks, they struggle in tasks requiring social norm and deep contextual awareness Hu et al. (2023); Sravanthi et al. (2024).
@@ 5.2 Social Dynamics
Politeness is vital in human interaction and has been studied in relation to conversational derailment Zhang et al. (2018b) and constructiveness De Kock and Vlachos (2021); Zhou et al. (2024). Computational approaches to politeness detection have evolved from ML-based classifiers using domain-independent markers Danescu-Niculescu-Mizil et al. (2013) to successfully leveraging LLMs for annotation Zhou et al. (2024); Ziems et al. (2024),
@@
Power and status influence conversational dynamics, affecting language use and turn-taking. Low-power individuals tend to mimic high-power speakers’ linguistic styles more than vice versa Danescu-Niculescu-Mizil et al. (2012). Hence, higher status roles can control the flow of discussions, foster social inequalities and hence degrade discussion quality. LLMs perform well in identifying power differentials in discussions Ziems et al. (2024).
@@
Disagreements, when constructive, enhance discussions by fostering deeper understanding Friess (2018); De Kock and Vlachos (2021). Measuring disagreement levels is complex, with frameworks such as Graham’s hierarchy (ranging from name calling to refuting the central point; Graham, 2008) and dispute tactics Walker et al. (2012); Benesch et al. (2016); De Kock et al. (2022) providing structured analyses. LLMs have been successfully employed as dispute tactics annotators Zhou et al. (2024).
@@ 5.3 Emotion and Behavior
Empathy fosters constructive discussions and is commonly measured as Perceived Empathy Concannon et al. (2023). Approaches range from coding schemes to linguistic markers Macagno et al. (2022). Evaluations of LLMs in empathy detection tasks show that empathy understanding is challenging for LLMs Ziems et al. (2024); Xu and Jiang (2024).
@@
Toxicity in online discussions refers to harmful or disrespectful language that hinders productive discourse Avalle et al. (2024). Identifying toxicity is vital to maintain constructive communication. LLMs have been proven to be adept at identifying toxicity Ziems et al. (2024).
@@
Sentiment analysis gauges the emotional tone of discussions, which influences the quality of interactions De Kock and Vlachos (2021). It helps identify whether discussions are positive, negative, or neutral. GPT-4 has successfully been utilized as a sentiment annotator Zhou et al. (2024). Ziems et al. (2024) also evaluated LLMs in figurative language identification, including sarcasm, revealing moderate performance and occasional misclassification.
@@
Controversy arises from divergent viewpoints, leading to polarized exchanges. The spread of political leanings among discussion participants of controversial topics and sentiment distribution analysis are common approaches to measure it Avalle et al. (2024). Ziems et al. (2024) evaluated GPT-4 in political ideology identification, showing moderate predictive performance.
@@
Constructiveness fosters meaningful dialogue, especially in online discussions, by promoting resolution and cooperation Shahid et al. (2024). It is signalled by linguistic features De Kock et al. (2022); Falk et al. (2024) and empirical research develops models to predict it Zhou et al. (2024). Shahid et al. (2024) found that GPT-4 preferred dialectical over logical arguments in assessing constructiveness. They also found that human participants rated LLM-generated and human-AI co-written comments as significantly more constructive than those written independently by humans.

### v2

@@ Abstract
We present a survey of methods for assessing and enhancing the quality of online discussions, focusing on the potential of Large Language Models (LLMs). While online discourses aim, at least in theory, to foster mutual understanding, they often devolve into harmful exchanges, such as hate speech, threatening social cohesion and democratic values. Recent advancements in LLMs enable artificial facilitation agents to not only moderate content, but also actively improve the quality of interactions. Our survey synthesizes ideas from Natural Language Processing (NLP) and Social Sciences to provide (a) a new taxonomy on discussion quality evaluation, (b) an overview of intervention and facilitation strategies, (c) along with a new taxonomy of conversation facilitation datasets, (d) an Large Language Model (LLM)-oriented roadmap of good practices and future research directions, from technological and societal perspectives.
@@ 1 Introduction
Social Science provides theories and applications for the facilitation of a discussion, but in specific contexts, such as teaching/learning Mansour (2024) or clinical discussions Gelula (1997), without much research devoted to online discussions, such as in social media. While prior NLP studies have explored LLM-facilitated discussions Burton et al. (2024); Aher et al. (2023); Beck et al. (2024); Schroeder et al. (2024); Small et al. (2023); Cho et al. (2024), rarely does Social Science work examine how facilitation can be automated Gimpel et al. (2024).
Figure 1: A conceptualization of this survey. We explore approaches from different disciplines, which recommend their own ways of evaluating and improving discussions.
In this survey, we combine LLM-based methods, with ideas from Social Science (e.g., Deliberative Theory) when discussing how to evaluate online discussions, and when exploring intervention strategies. Figure 1 provides a high-level conceptualization of our work.
The main research question of this survey is can LLMs be used effectively as facilitators in online discussions? To explore this question, we focus on three key areas: (1) methods (potentially also LLM-based) for evaluating aspects of online discussions, (2) intervention strategies for facilitation, and (3) available data resources relevant to facilitation. Specifically, we survey discussion evaluation aspects and introduce a new taxonomy (§4). We map tasks suited for ML models, LLMs, and humans, aggregate multidimensional insights on facilitation strategies (§5), and outline future possibilities for LLMs6). Additionally, we compare major datasets, dividing them into categories per task (§7). Our work focuses mostly on written thread-like discussions (§2).
Our findings show that (a) many discussion evaluation dimensions coexist in the literature; (b) LLM advancements show significant promise in improving the quality and timeliness of facilitation methods; (c) while surveying the existing datasets, we notice a scarcity of datasets for studying facilitation. We posit that LLM-generated discussions, could become an asset to develop and test automatic facilitation strategies in diverse artificial discussions, before testing the strategies and the LLM-based facilitator agents in more costly experiments with human participants.
@@ 2 Terminology
Given the numerous aspects to consider regarding discussion quality and facilitation, we clarify the terminology we use. We highly recommend consulting the Terminology Section of Appendix C and, especially, Table 3, where we explain our findings with regard to the terms used in the literature.
@@
The term ‘moderation’ is more commonly used in NLP Argyle et al. (2023), typically referring to the flagging and/or removal of unwanted content (‘content moderation’), while ‘facilitation’ is more prevalent in the Social Sciences, where it encompasses a broader scope, including active interventions Vecchi et al. (2021); Kaner et al. (2007); Trenel (2009). Given the limited attention to facilitation in NLP and the survey’s grounding in Social Science, we distinguish between the terms, even though they are sometimes used interchangeably in the literature.
@@
This survey mainly focuses on ‘Real-Time, Ex-Post-moderation’, i.e., moderation happening just after the user has posted some content. This is different from pre-moderation approaches, such as nudging users before they post harmful content Argyle et al. (2023), or delaying the posting of user content until a moderator has had the chance to check it.
@@
The definitions of these terms often vary across literature Russmann and Lane (2016); Goñi (2024). We focus on discussions, a general term for verbal/written exchanges Russmann and Lane (2016), and deliberations, a term for structured discussions focusing on opinion sharing Degeling et al. (2015); Lo and McAvoy (2023). This is in contrast to the (at least in theory) collaborative nature of dialogues Rose-Redwood et al. (2018); Bawden (2021); Goñi (2024) and the competitive and organized nature of debates Lo and McAvoy (2023).
@@
(or “threads”) are discussions which start from an Original Post (OP) with subsequent comments replying to either the OP or to other comments Seering (2020).
@@ 3 Comparison to Other Surveys
Only two studies have surveyed the field of NLP while also considering ideas from Social Science. However, they focus mainly on Argument Mining (AM). These are the studies of Wachsmuth et al. (2024) and Vecchi et al. (2021). Wachsmuth et al. (2024) focus primarily on discussion evaluation disregarding its relation to facilitation, which is one of the main goals of our survey. The survey of Vecchi et al. (2021) argues that advancing AM for social good requires a collaborative effort between AM and Social Science. They point out that traditional AM has prioritized the logical structure and soundness of arguments, while overlooking other important dimensions, such as civility, respectfulness, inclusiveness, originality, and the broader impacts of discussions—such as encouraging mutual understanding and problem-solving. Building on these notions, we incorporate ideas from Social Science into NLP-based approaches, discussing both discussion evaluation and facilitation, both with a focus on the potential of LLMs.
@@ 4 Discussion Quality Evaluation
Improving online discussions presupposes being able to define and measure discussion quality. While there have been attempts to provide frameworks for discussion quality evaluation Kies (2022); Gerber et al. (2018), none of them is directed towards facilitation. Crucially, most existing frameworks ultimately rely on human judgments as their reference point, yet human evaluation is expensive, slow, and shows low inter-rater agreement on dimensions that involve subjective interpretation, such as pragmatic cues (Smith et al., 2022; Yeh et al., 2021; Khalid and Lee, 2022). This evaluation bottleneck motivates a taxonomy of evaluation methods that is both comprehensive and amenable to scalable automatic measurement.
In this work, we draw from the works of Bächtiger et al. (2022, 2010); Steenbergen et al. (2003); Falk and Lapesa (2023) and Kies (2022) to define a new social-science-informed taxonomy for discussion quality dimensions. While we present a structured taxonomy, it is important to note that the categories are not mutually exclusive. Rather, elements within the taxonomy may coexist within evaluation dimensions, complement one another, or serve as explanatory mechanisms for other dimensions. An example of the dimension interaction can be found in Table F in the Appendix. The grouped dimensions along with the NLP approaches are shown in the Appendix in Table 4.
@@ 4.1 Structure and Logic
Argument Quality (AQ) is a multidimensional concept assessed through logical, rhetorical, and dialectical dimensions Wachsmuth et al. (2017). The logical dimension focuses on the coherence and structure of the argument. The rhetorical dimension assesses persuasiveness, focusing on the argument’s style and emotional appeal. The dialectical dimension assesses the constructiveness of the argument. Empirically, threads with well-formed claim-evidence chains exhibit higher coherence and lower odds of devolving into ad-hominem attacks, making AQ scores, as a discussion quality dimension, an early-warning indicator of derailment (Chang and Danescu-Niculescu-Mizil, 2019). All the above dimensions of automatic argument-structure analysis can be used by a facilitator to keep the discussion fact-centered, inclusive, and on track Falk et al. (2021); Falk and Lapesa (2023).
@@
‘Coherence’, as described above, evaluates logical consistency, while ‘flow’ assesses smooth progression in discussions Li et al. (2021). Both are essential tools for facilitators in their effort to redirect off-topic comments and guide transitions between topics during a discussion Lambert et al. (2024); Park et al. (2012); Falk et al. (2024). A sudden drop in how well responses match the topic or question often comes before personal attacks or off-topic turns (Chang and Danescu-Niculescu-Mizil, 2019; Zhang et al., 2018), making coherence and flow indicators of argument structure and a valuable early signal for facilitators.
@@
How speakers alternate, the frequency of their turns, and the participants they address can serve as a diagnostic of conversational health. Balanced exchanges enhance coherence (Cervone and Riccardi, 2020), predict constructiveness (§4.3) (Niculae and Danescu-Niculescu-Mizil, 2016), and provide facilitators with actionable cues (Schroeder et al., 2024). To gauge speaking time, turn count, and word usage, researchers have applied metrics such as entropy Niculae and Danescu-Niculescu-Mizil (2016) and Gini coefficients Schroeder et al. (2024).
Linguistic Markers
Linguistic markers have been used to help model content and expression in online discussions Wilson et al. (1984). Early methods used lexicons for sentiment, toxicity, politeness (§4.2 and 4.3) and collaboration evaluation Lawrence et al. (2017); Avalle et al. (2024). For example, spikes in hedges (e.g., ‘maybe’, ‘I guess’) invite clarification requests by facilitators, while bursts of second-person pronouns, similarly to turn-taking, often foreshadow personal attacks and can prompt a civility nudge Niculae and Danescu-Niculescu-Mizil (2016).
@@
Rooted in Speech Act Theory Austin (1975); Searle (1969), dialogue acts have been employed to assess deliberative quality and analyze facilitation strategies Fournier-Tombs and MacKenzie (2021); Chen et al. (2024a). They characterize dialogue turns (e.g., interruption) to analyze interaction dynamics Ferschke et al. (2012); Stolcke et al. (2000); Zhang et al. (2017); Al-Khatib et al. (2018). Positive (e.g., causal reasoning) or negative (e.g., disrespect) dialogue acts can be scored to reflect discussion quality and low scores may indicate the need for interventions Ziems et al. (2024); Cimino et al. (2024); Martinenghi et al. (2024); Schroeder et al. (2024).
@@
Pragmatic comprehension—how context shapes meaning—is crucial to facilitation, as intended meanings often diverge from literal expressions (i.e., implicature). Humans resolve such ambiguity using social and commonsense knowledge. Grice’s maxims Grice (1975), a central pragmatic concept, can help explain this process by outlining the conversational principles people rely on to infer meaning, while they have already been used to assess discussion quality Jwalapuram (2017); Langevin et al. (2021); Ngai et al. (2021); Nam et al. (2023).
@@ 4.2 Social Dynamics
Politeness serves as a cornerstone of prosocial behavior, an attribute that facilitators desire to foster in online discussion forums Lambert et al. (2024). In the context of facilitation, it has mainly been studied in relation to conversational derailment (§7) Zhang et al. (2018) and constructiveness (§4.3) De Kock and Vlachos (2021); Zhou et al. (2024).
@@
Power and status influence conversational dynamics, affecting language use and turn-taking (§4.1). Higher status speakers can control the flow of discussions and foster social inequalities. Interestingly, low-status individuals tend to mimic the linguistic styles of high-status speakers more than the opposite Danescu-Niculescu-Mizil et al. (2012), and this can be used as a signal that there is high/low-status imparity in a discussion. Facilitators may intervene, then, to ensure that the right to speak is evenly distributed among participants, preventing projection of social biases and stereotypes.
@@
Disagreements, when constructive, improve discussions by fostering deeper understanding Friess (2018); De Kock and Vlachos (2021). Assessing disagreement, however, is complex. The hierarchy of Graham, 2008 considers disagreement tactics ranging from name calling to refuting the central point. Along with other work on dispute tactics Walker et al. (2012); Benesch et al. (2016); De Kock et al. (2022), it can be used to examine types of disagreements in a discussion.
@@ 4.3 Emotion and Behavior
Empathy is the ability to understand others’ perspectives and emotions and respond correspondingly Lipman (2003); Xu and Jiang (2024). Facilitators desire to foster empathy in online discussions, since it encourages prosocial behavior and boosts engagement Xu and Jiang (2024); Concannon and Tomalin (2024); Lambert et al. (2024). To do so, they encourage users to share personal stories and experiences Schroeder et al. (2024). Various coding schemes Macagno et al. (2022), psychological indicators (e.g., the emotion-laden words of Furniturewala and Jaidka, 2024), and dimensions (e.g., perceived engagement such in Xu and Jiang, 2024) have been used to detect both expressed and perceived empathetic traits.
@@
Toxicity in online discussions refers to harmful or disrespectful language that hinders productive discourse and can derail meaningful discussions Avalle et al. (2024). Facilitation is key to maintaining healthy communication, requiring both early detection of toxicity and (in the case of more active facilitation) proactive de-escalation strategies, such as conversation redirection or positive engagement (§5). In the case of conventional moderation that only aims to flag or remove toxic content, debate persists over what content warrants removal Warner et al. (2025); Habibi et al. (2024); Pradel et al. (2024).
@@
Sentiment analysis helps identify whether discussions are positive, negative, or neutral. In the context of facilitation, sentiment analysis gauges the tone of discussions, which influences the quality of interactions De Kock and Vlachos (2021). Positive sentiment contributions in online discussion forums usually signal prosocial behavior and hence are highly encouraged by facilitators Lambert et al. (2024), while negative sentiments among discussants contribute to conversation toxicity Avalle et al. (2024).
@@
Controversy arises from divergent viewpoints, leading to polarized exchanges that can escalate to toxicity and derail online discussions Avalle et al. (2024). Controversial comments have been shown to contribute to a decline in positive emotions and a sustained rise in anger Hessel and Lee (2019); Chen et al. (2024b). The spread of political leanings among discussants and sentiment distribution analysis are common approaches to measure controversy Avalle et al. (2024).
@@
Constructiveness fosters meaningful dialogue, especially in online discussions, by promoting resolution and cooperation Shahid et al. (2024). It is often signalled by linguistic markers (§4.1) De Kock et al. (2022); Falk et al. (2024). A facilitator can exploit a constructiveness score; threads trending upward are worth highlighting or summarizing, whereas a downward drift may trigger facilitation tactics such as slower, structured turn-taking or clarification prompts (De Kock and Vlachos, 2021).